OpenStudy (anonymous):

Find the Maclaurin expansion of cos(x^3)

OpenStudy (anonymous):

So i'm thinking since it's Maclaurin a = 0. I know the Maclaurin expansion of cos(x) = SUM(-1)^n * (x^(2n))/(2n)!

OpenStudy (anonymous):

replace x with x^3

OpenStudy (anonymous):

why would we just do that ? if it was asking for the maclaurin expansion of x cos (x) i would just multiply the expansion of cos(x) by x. So wouldn't I need to multiply it by something again or do some computation to the orignal expansion besides just replacing x ?

OpenStudy (anonymous):

just replacing

OpenStudy (anonymous):

for example u know that \( \frac{1}{1-x}=1+x+x^2+... \) for |x|<1 now for \( \frac{1}{1-x^2}\) u just need to replace x with x^2 to get \( \frac{1}{1-x^2}=1+x^2+x^4+... \)

OpenStudy (anonymous):

oh wow. I get it now. sorry for making you have to type up those equations. Thanks though!

OpenStudy (anonymous):

welcome my friend